OpenStudy (anonymous):

HELP ! Brandon opened a credit card at a department store with an APR of 14.99% compounded monthly. What is the APY on this credit card? 1.25% 29.98% 19.55% 16.06%

OpenStudy (amistre64):

\[(1+\frac{apr}{12})^{12}=1+apy\]

OpenStudy (amistre64):

APR is a compounding interest APY is the simple interest that would have amounted to the same ....

OpenStudy (anonymous):

(1+14.99%/12)1+14.99 <<< ? Did I set it up right ?

OpenStudy (amistre64):

almost, lets convert the % to its decimal .... (1+.1499/12)^(12) = 1 + apy

OpenStudy (anonymous):

(1+0.1499/12)1+apy ?

OpenStudy (amistre64):

thats looking better, but your posts are not actually formatting the same as mine .... (1+0.1499/12)^(12) = 1+apy \[(1+\frac{0.1499}{12})^{12} = 1+apy \]

OpenStudy (anonymous):

oh ! :D what is apy ?

OpenStudy (amistre64):

"What is the APY on this credit card?" apy is the ... well ... apy that you are trying to determine. A nnual P ercentage Y ield

OpenStudy (anonymous):

I got this (1.0124916667)^12=1+apy

OpenStudy (amistre64):

so far so good

OpenStudy (anonymous):

im stuck on (1.0124916667)^12=1+apy I dont know what else to do :c

OpenStudy (amistre64):

work up the power of 12 ...

OpenStudy (amistre64):

(1.0124916667)^12 your calculator should have a "^" key, and some "(", ")" keys

OpenStudy (amistre64):

if its the windows basic calculator, there should an \(X^y\) option

OpenStudy (anonymous):

Its gave me 1.6063988....

OpenStudy (amistre64):

thats incorrect, you prolly hit a wrong key

OpenStudy (amistre64):

(1.01249)^12 = 1.1606169 .....

OpenStudy (amistre64):

1.1606169 = 1 + apy subtract 1 from each side and convert to %

OpenStudy (anonymous):

I GOT ITTT !! :D 16.06%

OpenStudy (amistre64):

correct :)

OpenStudy (anonymous):

Thank you soo much ! I really was lost !

OpenStudy (amistre64):

you did fine, good luck ;)